OpenStudy (anonymous):

The area of a square piece of cardboard is 81 square inches. What is the length of the cardboard?

OpenStudy (anonymous):

9 inches because x^2=81 and x=9 inches :)

OpenStudy (apolloschariot):

square has equal sides and area is length*breadth or u can say that area of square is (side)^2 let side be L area is 81 81=(L)^2 9=L